CEF5CONV(1) CEF5CONV(1) NAME cef5conv−convertaTeXdocumentinBig5encodingcontainingCEFmacrosinto‘preprocessed’ form. SYNOPSIS cef5conv<infile>outfile DESCRIPTION TheChineseEncodingFramework(CEF)developedbyChristianWittern <cwittern@conline.central.de>makesitpossibletoincludeCJKcharactersinvariousencodingsusingtheSGMLmacros&CX-aabb;forBig5andCNSencodingand&U-aabb;forUni-code(Xdenotesanumberbetween0and7;0representsBig5encoding,1-7CNSencoding planes1-7.aabbiseitheradecimalorhexadecimalnumber). cef5convconvertstheCEFmacrosofaBig5encodeddocumenttogetherwiththeBig5encoding itselfintoaformwhichcandirectlyprocessedwithLATEX2ε.Itisidenticalto bg5conv<infile|cefconv>outfile ThisfilterispartoftheCJKmacropackageforLATEX2ε. SEEALSO sjisconv(1),bg5conv(1),cefconv(1),cefsconv(1),extconv(1), theCJKdocumentationfiles. AUTHOR WernerLemberg<wl@gnu.org> CJKVersion4.3.0 20-Jun-1999 1